OpenStudy (anonymous):

simplify (x^3/5)^7/6. no clue where to begin on solving

OpenStudy (anonymous):

(a^x)^y = a^(xy) also, (m/n) *(p/q) = mp/nq

OpenStudy (anonymous):

so what numbers plug into where. I definetly missed this part in class

OpenStudy (anonymous):

no, you dont plug anything anywhere. I was just showing rules of how exponents and multiplying fractions work.

OpenStudy (anonymous):

(x^3/5)^7/6 = x^(3/5)*(7/6) so since (3/5)*(7/6) = (21/30) = (7/10) the answer is x^(7/10)

OpenStudy (anonymous):

the rules is " a raised to the power x raised to the power y is equal to a raised to the power product of x and y"

OpenStudy (anonymous):

(x^3/5)^7/6 = x^(3/5)*(7/6) so since (3/5)*(7/6) = (21/30) = (7/10) the answer is x^(7/10)

OpenStudy (anonymous):

so say if it was switched to (x^5/3)^9/4 it would be 45/12 = 15/4

OpenStudy (anonymous):

yes that would be correct x^(15/4)

OpenStudy (anonymous):

yes, grags, very good!

OpenStudy (anonymous):

to check this...plug a number into your calculator for x. lets say x=5, so (5^5/3)^9/4 = 417.9626906 now plug 5 into what you think is the correct answer of x^(15/4) 5^(15/4) also equals 417.9626906 so your answer checks

OpenStudy (anonymous):

ok lets hope i remember that. Final is tomorrow and i am doing a final review. somehow after 40 hours this week of studying i still am getting hung up on some questions